OpenStudy (anonymous):

Which of the following disjunctions is false? 2 · 3 = 6 or 4 + 5 = 10 5 - 3 = 2 or 3 · 4 = 12 3 + 4 = 9 or 5 · 2 = 11 6 · 2 = 11 or 3 + 5 = 8

OpenStudy (sirm3d):

the third disjuction is false. both statements are false

OpenStudy (anonymous):

1. 4+5= 9 not 10 2. both are right 3. both are false 4. 6*2 = 12 not 11

OpenStudy (anonymous):

you got it??

OpenStudy (anonymous):

ye i think so

OpenStudy (anonymous):

good

OpenStudy (anonymous):

1)4+5=10 3)3+4=7 5.2=10 4)6.2=12